Quiz I4. Monitor a phone call. Classify the call as a voice call (V), if someone is speaking: or a data call (D) if the call is carrying a modem or fax signal. Classify the call as long (L) if if the call lasts for more than three minutes; otherwise classify the call as brief (B). Based on data collected by the telephone company, we use the following probability model. PV-07, P14-0.6, PIVL 0.35. Please find the following probabilities: (a) PDL (c) PIVB (e) PIVUD (b) PDUL (d) PVUL (D PILB
Answer #1

(a)

P(DL) = P(L) - P(VL) = 0.6 - 0.35 = 0.25

(b)

P(D) 1 P(V) = 1 - 0.7 = 0.3

So,

P(DUL) P(D) P(L)- P(Dn L) = 0.3 +0.6- 0.25 0.65

(c)

P(VB) = P(V) - P(VL) = 0.7 - 0.35 = 0.35

(d)

P(V UL)= P(V)+ P(L) - P(Vn L) = 0.7+ 0.6- 0.35 0.95

(e)

P(V U D) = P(V)+ P(D) = 1

(f)

Since events L and B are mutually exclusive so

P(LB) =P(L and B)= 0